how much money would be in an account after 5 years if an original investment of $6500 was compounded quarterly at 4.5%?. compare this amount to the same investment that was compounded daily. round to the nearest cent.
Found 2 solutions by solver91311, lwsshak3:
Answer by solver91311(24713)





Where is the future value, is the initial investment, is the interest rate expressed as a decimal, is the number of compounding periods per year (4 for quarterly, 365 for daily), and is the number of years.

Get out your calculator.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by lwsshak3(11628)

how much money would be in an account after 5 years if an original investment of $6500 was compounded quarterly at 4.5%?. compare this amount to the same investment that was compounded daily. round to the nearest cent.
**
Compound Interest formula:
A=P(1+i)^n, P=initial investment, i=interest rate per period, A=amount after n periods
For given problem:
compounding quarterly
P=6500
i=.045/4 (annual interest/number of compounding periods per year)
n=5*4=20 periods
A=6500(1+.045/4)^20
using calculator
A=8129.88
..
compounding daily
P=6500
i=.045/365 (annual interest/number of compounding periods per year)
n=5*365=1825 periods
A=6500(1+.045/365)^1825
using calculator
A=8139.98
..
compounding daily gives about $10 more over a 5-year period